# FlippedRAG: Black-Box Opinion Manipulation Adversarial Attacks to Retrieval-Augmented Generation Models

### v3

@@ 4.4. Baselines
To the best of our knowledge, no existing studies have specifically addressed opinion manipulation in black-box RAG systems. We therefore selected black-box PoisonedRAG and GARAG as baselines for opinion manipulation. Furthermore, we extended the baseline selection to include attack methods which target either LLMs or retrieval models and can be applied to black-box attack.
(1) Prompt Injection Attack. Prompt injection attacks constitute a class of adversarial exploits wherein malicious actors strategically insert compromised instructions or corrupted data into the inputs of LLM-integrated systems, thereby coercing the model’s outputs to align with predefined adversarial objectives(Liu et al., 2024).
(2) Disinformation. It (Pan et al., 2023) is a methodology for generating credible-sounding misinformation by leveraging LLMs, employing the adversarial prompt to induce LLMs to produce context passages imbued with specific opinions to the controversial topics.
(3) Static Text. We add static text to some documents that contradict the target opinion (“Experts dispute the above claims as being unreliable”), thereby inducing the LLM to reinforce the target opinion.
(4) PAT Transfer-based Attack. We apply retrieval adversarial strategy PAT(Liu et al., 2023b) to the surrogate model that has not undergone black-box imitation and transfer adversarial triggers to the RAG system, assessing the effectiveness of the black-box imitation.
(5) GARAG. It (Cho et al., 2024) employs genetic algorithms to optimize the discovery of novel adversarial documents that achieve dual objectives: maintaining retrievability within RAG systems, and deviating LLM outputs from factual accuracy.
(6) PoisonedRAG. It (Zou et al., 2024) is black-box attack framework leverages query insertion and LLM generation to construct malicious documents. Detailed baseline introduction is in Appendix A.2.
@@ 4.5. Evaluation Metrics
For RQ1, we aim to evaluate the relevance ranking ability of the surrogate model trained through black-box imitation of $\text{RAG}_{\text{black}}$ on standard ranking tasks and compared with the target retrieval model. Additionally, when adversarial triggers are applied to the target retrieval model, we compare the manipulation effect of triggers generated from the surrogate model versus those generated from the original BERT model, which is not trained by black-box imitation, to assess the effect of black-box imitation. We use the following evaluation metrics, higher values of these metrics demonstrate enhanced efficacy in black-box imitation, indicating that the surrogate model successfully acquires the ranking capabilities of the target model.
Normalized Discounted Cumulative Gain (NDCG). NDCG is designed to evaluate the effectiveness of search engines in ordering results by relevance.
@@
For RQ2, we aim to evaluate the manipulation performance of FlippedRAG on retriever rankings and LLM responses. We use Top3v, Top6v, Ranking Attack Success Rate (RASR)and Boost Rank (BRank) to reflect how effective our attack distort the retrieval ranking of RAG. Elevated values of these metrics demonstrate enhanced manipulation efficacy in ranking and generated opinion manipulation.
Top3v and Top6v. Top3v quantifies the enhancement in the proportion of target opinion within the top-3 rankings after manipulation. Top6v follows the same definition.
Ranking Attack Success Rate (RASR). RASR measures the average rate of successfully boosted candidates in the ranking of each query.
@@
We also define Opinion Manipulation Success Rate (OMSR) and Average Stance Variation (ASV) to evaluate the significance of our opinion manipulation on the LLM responses.
@@ 5. Experimental Results Analysis
Table 2. Comparison(%) of ranking ability between the surrogate model and the target retrieval model.
@@
As shown in Table 2, the surrogate model $M_{i}$ trained by black-box imitation is similar to the target retrieval model in relevance ranking ability for their close NDCG@10 scores. Their ranking results are also similar enough, validating the effectiveness of the black-box imitation. For convenience, we use $M_{Contriever}$ to represent the surrogate model that has its training data collected from $\text{RAG}_{\text{black}}$ which uses Contriever. The method of representation applies to surrogate models based on ranking data of Co-Condenser and ANCE.
@@
Table 3. Comparison of manipulation effect between the surrogate model $M_{i}$ and the BRET model without imitation.
Target Retrieval Model Target Opinion Victim Model Ranking Manipulation Performance Opinion Manipulation on Llama-3
$Top3_{v}$ $Top6_{v}$ $BRank$ $RASR\%$ $OMSR\%$ $ASV$
Contriever PRO w/o imitation 0.31 0.28 72.33 70.73 43.33 0.50
imitation 0.37 0.38 111.03 74.22 53.33 0.67
CON w/o imitation 0.27 0.24 67.1 71.93 40.00 0.27
imitation 0.37 0.35 105.13 78.26 53.33 0.70
Co-Condenser PRO w/o imitation 0.22 0.12 43.53 58.91 23.33 0.20
imitation 0.27 0.22 63.70 62.01 43.33 0.27
CON w/o imitation 0.07 0.13 37.43 57.11 20.00 0.07
imitation 0.17 0.19 58.96 64.42 33.33 0.27
ANCE PRO w/o imitation 0.30 0.22 57.20 62.79 43.33 0.43
imitation 0.38 0.32 85.96 69.76 53.33 0.47
CON w/o imitation 0.24 0.19 50.90 64.42 30.00 0.20
imitation 0.33 0.32 89.40 74.50 36.66 0.37
@@
We place great importance on the manipulation performance of the adversarial triggers generated from the surrogate model $M_{i}$. The ranking manipulation effect of the surrogate model $M_{i}$ on the target retrieval model should fall between the manipulation effect of the white-box target retrieval model and that of BERT-based transfer attacks without imitation as $M_{i}$ has acquired some of the target model’s capabilities. Only if triggers from the surrogate model has the better ranking manipulation effect on the target retrieval model than that from the origin model without imitation, we consider the black-box imitation process to be successful.
@@
As shown in Table 3, we find that in terms of ranking manipulation effectiveness, the surrogate models after black-box imitation outperform the baseline models. In both RASR and BRank of target documents, triggers generated from the surrogate model achieve better transfer attack effects on the target retrieval model. Specifically, $M_{Contriever}$ has the best ranking manipulation performance on the triggers generated for transfer attacking. $M_{ANCE}$ demonstrates ranking manipulation effectiveness comparable to that of $M_{Contriever}$, whereas the manipulation effect achieved by $M_{CoCondenser}$ is less ideal by comparison. It indicates that Co-Condenser is more robust and reliable than Contriver and ANCE.
@@
Additionally, the surrogate models promote more documents with target opinion into top 3 range, thereby having a significant impact on the response of $\text{RAG}_{\text{black}}$. Thus, we can see that the transfer attack based on the surrogate model achieves stronger opinion manipulation by successfully flipping opinions of RAG response. Black-box imitation enables the surrogate model to learn the features of the black-box retrieval model, allowing triggers generated by the surrogate model to induce a greater shift in opinion polarity in LLMs’ response compared to the baseline model. Table 3 also indicates that adversarial triggers generated by $M_{Contriever}$ and $M_{ANCE}$ place more target documents in top 3 range, thus achieving relatively better performance on OMSR and ASV.
@@
In short, the above experiment indicates that the black-box imitation process is able to help BERT learn internal knowledge about the target black-box retrieval model.
@@ 5.2. RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
After the black-box imitation training, we conduct opinion manipulation experiments across different $\text{RAG}_{\text{black}}$ (based on Llama3, Mixtral, or Vicuna) with FlippedRAG. We also conducted a comparative evaluation of FlippedRAG against established baselines to assess its relative efficacy in executing opinion manipulation on black-box RAG architectures.
Table 5. Opinion fluctuations in RAG responses based on clean collections of different LLMs.
LLM OMSR(%) ASV
Llama-3 6.7 -0.03
Mixtral 10.0 0.07
Vicuna 6.7 0.06
As shown in Table 4, FlippedRAG achieves significant opinion manipulation effect on all three LLMs. As a whole, it has a 40% – 50% chance of successful opinion flipping and realizes about 0.46 in opinion polarity shift of range 0 to 2. Moreover, it reveals that FlippedRAG based on $M_{Contriever}$ and $M_{ANCE}$ are more effective in manipulating the RAG opinion with relatively higher OMSR and ASV value. That is because these attacks exhibit superior performance on ranking distortion.
We also observe that the impact of FlippedRAG varies in opinion manipulation effectiveness across different LLMs. Llama3 and Mixtral have higher OMSR score in contrast to Vicuna for their greater instruction-following ability. Thus, Llama3 and Mixtral are more susceptible to the adversarial documents. On the contrary, for ASV, Vicuna actually demonstrates a relatively greater shift in opinion polarity. We analyze that this is due to the better diversity in responses from Llama3 and Mixtral when answering questions, resulting in less stable opinion polarity in the their responses.
In some cases, opinion manipulation is unsuccessful, instead of turning the opinion of $\text{RAG}_{\text{black}}$ response to the target opinion or making it remain unchanged, manipulation causes the opinion polarity to shift in the opposite direction of the target opinion, thereby lowering the ASV value. The cause of these reverse manipulation results may also stem from the randomness of LLMs generation, although this influence is relatively minor.
@@
Figure 4. FlippedRAG’s overall effect of opinion manipulation of on LLama-3 with retrieval model Contriever.
[…]
From the perspective of overall opinion manipulation of $\text{RAG}_{\text{black}}$ response, as shown in Figure 4, FlippedRAG is able to significantly alter the opinion of RAG-generated content in a specified stance direction. When the target opinion is “Pro”, compared to the opinion distribution before manipulation, the proportion of responses holding a supportive stance (in gray) has significantly increased in the opinion distribution after manipulation. The converse holds true when the target opinion is specified as ’Con’. Graphical representations of the opinion manipulation effects on Mixtral and Vicuna are provided in the Appendix B.2.
Furthermore, FlippedRAG exhibits no adverse effects on the generation performance of non-target queries, as the target questions and normal questions remain orthogonal in their semantic feature space.
Figure 8 in Appendix B.1 presents a case study of FlippedRAG.
To further analyze the opinion manipulation results on RAG response at a finer granularity, we conduct experiments across diverse domain of controversial topics to evaluate its efficacy in cross-domain opinion manipulation. We focus on “Health”, “Society”, “Government”, “Education” four domains. FlippedRAG significantly increases the proportion of candidate items with target opinion in the Top 3 of the retrieval list, thereby guiding the LLM to change its opinion in the response. However, the manipulation effect varies across different themes. The detailed discussion is in Appendix B.3.
Table 6. Ranking manipulation and opinion manipulation results across attack baselines on target opinion Pro/Con .
@@
The comparative experiment results indicate that both PoisonedRAG and FlippedRAG demonstrate superior efficacy in the opinion manipulation task, with PoisonedRAG exhibiting marginally enhanced performance compared to FlippedRAG. While Prompt Injection Attack, Disinformation, Static Text, and PAT Transfer-based attack achieve moderate effectiveness in opinion manipulation, GARAG manifests the most suboptimal performance within this adversarial task paradigm.
Table 7. The controlled user experimental results on opinion manipulation in controversial topics.
Subject Opinion Polarity of User
Topic 1 Topic 2
A Group (before manipulation) 5.8 4.8
B Group (after manipulation) 6.4 8.2
@@
The superior performance of PoisonedRAG and FlippedRAG stems from their targeted optimizations of critical components in the RAG workflow: retrieval and generation. In the generation phase, PoisonedRAG leverages LLMs to synthesize content engineered to steer target systems, while FlippedRAG adopts a minimal data perturbation strategy by curating opinion-specific documents that guide RAG systems to extrapolate desired stances through summarization. PoisonedRAG’s enhanced efficacy derives from its retrieval optimization: injecting the question into documents, which maximizes retrieval probability by exploiting the semantic self-similarity. In contrast, although FlippedRAG improves adversarial document rankings through black-box retriever imitation, its effectiveness remains suboptimal compared to exact question-question matching paradigm.
@@
The diminished efficacy of the PAT Transfer-based method compared to FlippedRAG can be attributed to its lack of a black-box retriever imitation process. Both Prompt Injection Attack and Static Text exclusively target the generation phase while neglecting the retrieval in RAG architectures, leading to significant performance degradation when migrating LLM-targeted attack strategies to RAG scenarios. Although Disinformation fundamentally remains a generation-optimized attack method, its approach of constructing biased content based on the question nevertheless achieves partial retrieval prioritization enhancements.
@@
The limited adaptability of GARAG to black-box opinion manipulation tasks arises from two fundamental constraints. First, its methodology—originally designed for closed-domain QA to deviate RAG systems from factual correctness—exploits a vast generative space where any incorrect output suffices as a successful attack. In contrast, our opinion manipulation framework strictly confines LLM outputs to binary stance options (support/oppose), significantly increasing the attack complexity. Second, our implementation reveals white-box dependencies in GARAG’s genetic algorithm, which leverages internal retriever and LLM calculated values during adversarial document optimization. These architectural dependencies undermine its efficacy in more authentic black-box scenarios.

### v4

@@ 4.4. Baselines
To the best of our knowledge, opinion manipulation in black-box RAG has not been directly studied. We therefore use PoisonedRAG and GARAG as baselines and further include black-box–applicable attacks targeting either LLMs or retrieval models. A detailed introduction of the baselines is in Appendix A.2.
(1) Prompt Injection Attack. Prompt injection attacks constitute a class of adversarial exploits wherein malicious actors strategically insert compromised instructions or corrupted data into the inputs of LLM-integrated systems, thereby coercing the model’s outputs to align with predefined adversarial objectives(liu2024formalizing).
(2) Disinformation. It (pan2023risk) is a methodology for generating credible-sounding misinformation by leveraging LLMs, employing the adversarial prompt to induce LLMs to produce context passages imbued with specific opinions to the controversial topics.
(3) Static Text. We add static text to some documents that contradict the target opinion (“Experts dispute the above claims as being unreliable”), inducing the LLM to reinforce the target opinion. Additionally, we also introduce a stronger baseline by integrating Disinformation and Static Text.
(4) PAT Transfer-based Attack. We apply retrieval adversarial strategy PAT(liu_order-disorder_2023) to the surrogate model that has not undergone black-box imitation and transfer adversarial triggers to the RAG system, assessing the effectiveness of the black-box imitation.
(5) GARAG. It (cho2024typos) employs genetic algorithms to optimize the discovery of novel adversarial documents that achieve dual objectives: maintaining retrievability within RAG systems, and deviating LLM outputs from factual accuracy.
(6) PoisonedRAG. It (zou2024poisonedrag) is black-box attack leverages query insertion and LLM generation to construct malicious documents.
@@ 4.5. Evaluation Metrics
For RQ1, we aim to evaluate the relevance ranking ability of the surrogate model trained through black-box imitation of $\text{RAG}_{\text{black}}$ on standard ranking tasks and compared with the target retrieval model. Additionally, when adversarial triggers are applied to the target retrieval model, we compare the manipulation effect of the triggers generated from the surrogate model versus those generated from the original BERT model, which is not trained by black-box imitation, to assess the effect of black-box imitation. We use the following evaluation metrics, higher values of these metrics demonstrate enhanced efficacy in black-box imitation, indicating that the surrogate model successfully acquires the ranking capabilities of the target model.
Normalized Discounted Cumulative Gain (NDCG). NDCG is designed to evaluate the effectiveness of search engines in ordering results by relevance. NDCG@10 denotes the NDCG metric calculated over the top-10 ranked documents, where higher values signify superior ranking effectiveness.
@@
For RQ2, we aim to evaluate the manipulation performance of FlippedRAG on retriever rankings and LLM responses. We use Top3v, Ranking Attack Success Rate (RASR) and Boost Rank (BRank) to reflect how effective our attack distort the retrieval ranking of RAG. Elevated values of these metrics demonstrate enhanced manipulation efficacy in ranking.
Top3v. Top3v quantifies the enhancement in the proportion of target opinion within the top-3 rankings after manipulation.
Ranking Attack Success Rate (RASR). It is the average rate of candidates whose rankings are successfully boosted for each query.
@@
We also define Opinion Manipulation Success Rate (OMSR) and Average Stance Variation (ASV) to evaluate the significance of our opinion manipulation on the LLM responses. Higher values of these metrics demonstrate stronger efficacy of opinion manipulation.
@@ 5. Experimental Results Analysis
Table 2. Comparison(%) of ranking ability between the surrogate model after imitation training and the target retrieval model.
@@
As shown in Table 2, the surrogate model $M_{i}$ trained by black-box imitation is similar to the target retrieval model in relevance ranking ability for their close NDCG@10 scores. Their ranking results are also similar enough, validating the effectiveness of the black-box imitation. For convenience, we use $M_{Contriever}$ to represent the surrogate model that has its training data collected from $\text{RAG}_{\text{black}}$ which uses Contriever. The method of representation applies to surrogate models based on the ranking data of Co-Condenser, ANCE and Qwen3-Embedding-4b.
@@
Table 3. Comparison of manipulation effect between the surrogate model $M_{i}$ with and without (w/o) imitation.
Target Retrieval Model Target Opinion Victim Model Ranking Manipulation
Opinion Manipulation
on Llama-3
$Top3_{v}$
BRankBRank
$RASR\%$ $OMSR\%$ ASVASV
Contriever PRO w/o imitation 0.31 72.33 70.73 43.33 0.50
imitation 0.37 111.03 74.22 53.33 0.67
CON w/o imitation 0.27 67.1 71.93 40.00 0.27
imitation 0.37 105.13 78.26 53.33 0.70
Co-Condenser PRO w/o imitation 0.22 43.53 58.91 23.33 0.20
imitation 0.27 63.70 62.01 43.33 0.30
CON w/o imitation 0.07 37.43 57.11 20.00 0.07
imitation 0.17 58.96 64.42 33.33 0.27
ANCE PRO w/o imitation 0.30 57.20 62.79 43.33 0.43
imitation 0.38 85.96 69.76 53.33 0.47
CON w/o imitation 0.24 50.90 64.42 30.00 0.20
imitation 0.33 89.40 74.50 36.66 0.37
Qwen3-Embedding-4b PRO w/o imitation 0.05 26.96 52.51 23.33 0.10
imitation 0.11 34.73 54.84 30.0 0.30
CON w/o imitation 0.13 30.43 55.50 23.33 0.13
imitation 0.16 39.90 58.10 40.0 0.33
@@
We place great importance on the manipulation performance of the adversarial triggers generated from the surrogate model $M_{i}$. The ranking manipulation effect of the surrogate model $M_{i}$ on the target retrieval model should fall between the manipulation effect of the white-box target retrieval model and that of BERT-based transfer attacks without imitation, since $M_{i}$ has acquired some of the target model’s capabilities. Only if triggers from the surrogate model have the better ranking manipulation effect on the target retrieval model than that from the origin model without imitation, we consider the black-box imitation process to be successful.
@@
As shown in Table 3, we find that in terms of ranking manipulation effectiveness, the surrogate models after black-box imitation outperform the baseline models. In both RASR and BRank of the target documents, triggers generated from the surrogate model achieve better transfer attack effects on the target retrieval model. Specifically, $M_{Contriever}$ has the best ranking manipulation performance on the triggers generated for transfer attack. $M_{ANCE}$ demonstrates ranking manipulation effectiveness comparable to that of $M_{Contriever}$, whereas the manipulation effect achieved by $M_{CoCondenser}$ and $M_{Qwen3}$ is less ideal by comparison. It indicates that Co-Condenser and Qwen3-Embedding-4b are more robust and reliable than Contriver and ANCE. However, our experimental findings demonstrate that FlippedRAG maintains effective imitation and achieves significant ranking manipulation efficacy even against state-of-the-art embedding models like Qwen3-Embedding-4b, despite its superior ranking performance and enhanced robustness.
@@
Furthermore, the surrogate models promote more documents with target opinion in the top 3 range, thereby having a significant impact on the response of $\text{RAG}_{\text{black}}$. Thus, we can see that the transfer attack based on the surrogate model achieves stronger opinion manipulation by successfully flipping opinions of the RAG response. Black-box imitation enables the surrogate model to learn the features of the black-box retrieval model, allowing triggers generated by the surrogate model to induce a greater shift in opinion polarity in the LLMs’ response compared to the baseline model. Table 3 also indicates that adversarial triggers generated by $M_{Contriever}$ and $M_{ANCE}$ place more target documents in the top 3 range, thus achieving relatively better performance on OMSR and ASV.
@@
In short, the above experiment indicates that the black-box imitation process is able to help BERT learn internal knowledge about the target black-box retrieval model, even the state-of-the-art embedding model. Furthermore, FlippedRAG achieves effective manipulation across diverse retrieval models and LLMs, demonstrating significant robustness to architectural variations.
@@ 5.2. RQ2: How significant does opinion manipulation influence the response of the LLM in RAG?
After the black-box imitation training, we conduct opinion manipulation experiments across different $\text{RAG}_{\text{black}}$ (based on Llama-3, Mixtral, or Vicuna) with FlippedRAG. We also conducted a comparative evaluation of FlippedRAG against established baselines to assess its relative efficacy in executing opinion manipulation on black-box RAG architectures.
Table 5. Opinion fluctuations in RAG responses based on clean collections for different LLMs.
LLM Llama-3 Mixtral Vicuna
OMSR(%) 6.7 10.0 6.7
ASV -0.03 0.07 0.06
As shown in Table 4, FlippedRAG achieves a significant opinion manipulation effect on all three LLMs. As a whole, it has a 40% – 50% chance of successful opinion flipping and realizes about 0.46 in the opinion polarity shift of range 0 to 2. Moreover, it reveals that FlippedRAG based on $M_{Contriever}$ and $M_{ANCE}$ are more effective in manipulating the RAG opinion with relatively higher OMSR and ASV values. That is because these attacks exhibit superior performance on ranking distortion.
We also observe that the impact of FlippedRAG varies in opinion manipulation effectiveness across different LLMs. Llama-3 and Mixtral have higher OMSR scores in contrast to Vicuna due to their greater instruction-following ability. Thus, Llama-3 and Mixtral are more susceptible to the adversarial documents. On the contrary, for ASV, Vicuna actually demonstrates a relatively greater shift in opinion polarity. We analyze that this is due to the better diversity in responses from Llama-3 and Mixtral when answering questions, resulting in less stable opinion polarity in the their responses.
In some cases, opinion manipulation is unsuccessful, instead of turning the opinion of $\text{RAG}_{\text{black}}$ response to the target opinion or make it remain unchanged, manipulation causes the opinion polarity to shift in the opposite direction of the target opinion, thereby lowering the ASV value. The cause of these reverse manipulation results may also stem from the randomness of LLMs generation, although this influence is relatively minor.
@@
Figure 4. Overall opinion manipulation effect on Llama-3 with Contriever. We use color blocks to represent the distribution of different opinions and polarities corresponding to the generated content.:
From the perspective of general opinion manipulation of $\text{RAG}_{\text{black}}$ response, as shown in Figure 4, FlippedRAG is able to significantly alter the opinion of RAG-generated content in a specified stance direction. When the target opinion is “Pro”, compared to the opinion distribution before manipulation, the proportion of responses holding a supportive stance (in gray) has increased significantly in the opinion distribution after manipulation. The converse holds true when the target opinion is specified as ’Con’. Graphical representations of the opinion manipulation effects on Mixtral and Vicuna are provided in Figure 9 in Appendix B.2. Furthermore, FlippedRAG exhibits no adverse effects on the generation performance of non-target queries, as the target questions and normal questions remain orthogonal in their semantic feature space.
To further analyze the opinion manipulation results on the RAG response at a finer granularity, we conduct experiments across diverse domains of controversial topics to evaluate its efficacy in cross-domain opinion manipulation. We focus on “Health”, “Society”, “Government”, “Education” four domains. FlippedRAG significantly increases the proportion of candidate items with target opinion in the Top 3 of the retrieval list, guiding the LLM to change its opinion in the response. However, the manipulation effect varies across different themes. The detailed discussion is in Appendix B.3.
Table 6. Ranking manipulation and opinion manipulation results across attack baselines on target opinion Pro/Con . "-" denotes not applicable.
@@
The comparative experiment results indicate that both PoisonedRAG and FlippedRAG demonstrate superior efficacy in the opinion manipulation task, with PoisonedRAG exhibiting marginally enhanced performance compared to FlippedRAG. While Prompt Injection Attack, Disinformation, Static Text, "Disinformation + Static text", and PAT transfer-based attack achieve moderate effectiveness in opinion manipulation, GARAG manifests the most suboptimal performance within this adversarial task paradigm.
@@
The superior performance of PoisonedRAG and FlippedRAG stems from their targeted optimizations of critical components in the RAG workflow: retrieval and generation. In the generation phase, PoisonedRAG leverages LLMs to synthesize content engineered to steer target systems, while FlippedRAG adopts a minimal data perturbation strategy by curating opinion-specific documents that guide RAG systems to extrapolate desired stances through summarization. The enhanced efficacy of PoisonedRAG derives from its retrieval optimization: injecting the question into documents, which maximizes the retrieval probability by exploiting the semantic self-similarity. In contrast, although FlippedRAG improves adversarial document rankings through black-box retriever imitation, its effectiveness remains suboptimal compared to exact question-question matching paradigm.
@@
The diminished efficacy of the PAT transfer-based method compared to FlippedRAG can be attributed to its lack of a black-box retriever imitation process. Both Prompt Injection Attack and Static Text exclusively target the generation phase while neglecting the retrieval in RAG architectures, leading to significant performance degradation when migrating LLM-targeted attack strategies to RAG scenarios. Although Disinformation fundamentally remains a generation-optimized attack method, its approach of constructing biased content based on the question nevertheless achieves partial retrieval prioritization enhancements. The stronger baseline integrating Disinformation and Static Text demonstrates improved ranking relevance and guides LLM output during the generation phase, consequently achieving superior manipulation efficacy compared to its constituent baselines. However, due to the absence of dedicated optimization for retrieval or ranking relevance, its ranking manipulation capability remains constrained. Consequently, the overall attack efficacy falls short of both FlippedRAG and PoisonedRAG.
Table 7. The controlled user experimental results on opinion manipulation in controversial topics. Mean $\pm$ SD represents the arithmetic mean and standard deviation of user opinion polarity values. * denotes statistical significance at p < 0.05.
Topic Target Subject
Mean $\pm$ SD
Genetically Modified
Organisms
[…]
Corporate
Income Tax
[…]
Medical Aid
in Dying
[…]
@@
The limited adaptability of GARAG to black-box opinion manipulation tasks arises from two fundamental constraints. First, its methodology, originally designed for closed-domain QA to deviate RAG systems from factual correctness, exploits a vast generative space where any incorrect output suffices as a successful attack. In contrast, our opinion manipulation framework strictly confines LLM outputs to binary stance options (support/oppose), significantly increasing the attack complexity. Second, our implementation reveals white-box dependencies in GARAG’s genetic algorithm, which leverages internal retriever and LLM calculated values during adversarial document optimization. These architectural dependencies undermine its efficacy in more authentic black-box scenarios.